# Syntactic Language Change in English and German: Metrics, Parsers, and Convergences

### v1

@@ Results
We show the distribution of the texts identified as perfect sentences (without any issues), sentences with issues, and non-sentences in Figure 3. We observe that: (1) German data has more issues compared to English. For the older time periods, e.g., the first 3 decades in each corpus, 60-70% German sentences are found to contain issues, while there are only 20-30% English sentences that are found to be problematic. (2) Only a small portion of sentences were identified as non-sentences. Specifically, 6 out of 160 texts (3.75%) from DeuParl and 2 out of 220 texts (0.1%) from Hansard are found not to be sentences. (3) The proportion of problematic sentences is decreasing over time; for both languages, all sentences after 2000 are flawless.
@@
In summary, our annotations suggest that English data from Hansard is substantially less problematic than the German data from Deuparl. Notably, the most prevalent issues in the German data, namely historical spelling and OCR spelling errors, rarely appear in the English data. This leads us to question whether the Hansard Corpus has already undergone preprocessing like spelling normalization. Through contact with the UK Parliament enquiry services,888https://www.parliament.uk/site-information/contact-us/ we received feedback indicating that the data has not been preprocessed regarding spelling normalization. As for OCR errors, a statement on their website999https://hansard.parliament.uk/about explicitly mentions that OCR errors are curated once identified. As a consequence, we conclude that English has undergone smaller changes in terms of spelling in the past 200 years compared to German, making German the more interesting (and much less researched) language from a historical perspective.
@@ 3.2 Dataset Construction
We apply the preprocessing pipeline introduced above to get up to 200k sentences for each decade and language. To make the English and German data comparable, we limit the time period for English to the decade 1860-2020. Next, we tokenize the sentences and do part-of-speech tagging using Stanza (Qi et al., 2020). The length of sentences is then the number of tokens after tokenization. Figure 5 and 6 show the distribution of sentence lengths and the average sentence length over time. We see that shorter sentences appear more often in the later periods and thus the average sentence length is overall decreasing over time for both languages. For instance, the average sentence length in DeuParl decreases from 35 to 20 from the time window 1860-1870 to 2000-2020. The only exception is the last subtrend for Hansard, where the average length increases by 2 tokens. Metrics like mDD are highly sensitive to sentence length (Lei and Jockers, 2020), thus, it is important to control for sentence length when constructing the dataset for observation.
@@ 3.2.1 Final Dataset
We select sentences with varying number of tokens for observation: 5, 10, 15, 20, 30, 40, 50, 60, and 70, covering short, middle and long sentences. We aim to observe syntactic changes in sentences of the same length (i.e., same amount of tokens). Nevertheless, to obtain sufficient sentences for each length and time period, we set an offset of 2 tokens for sentence grouping, e.g., sentences of a length from 10 to 12 are grouped together; in addition, we consider decade groups with each spanning 20 years. For both English and German, the final dataset contains 450 sentences per length and decade group. These sentences are randomly sampled from a total of 32,400 preprocessed sentences for each language.
@@ 4 Dependency Parsers
Unlike previous approaches, which often solely relied on a single parser, typically the Stanford CoreNLP Parser (Manning et al., 2014), we will base our observations on various parsers with the two most popular design choices: transition-based and graph-based. Transition-based parsers sequentially predict the arc step by step based on the current state, while graph-based parsers globally optimize the dependency tree, aiming to find the highest-scored one. The parsers used in this work include: (1) transition-based: CoreNLP (Manning et al., 2014), StackPointer (Ma et al., 2018); (2) graph-based: Biaffine (Dozat and Manning, 2016), Stanza (Qi et al., 2020), TowerParse (Glavaš and Vulić, 2021) and CRF2O (Zhang, Li, and Zhang, 2020).
We train model checkpoints for StackPointer, Biaffine, and CRF2O on the concatenated train set of the treebanks in UD2.12, separately for each language.101010We use the implementation of Biaffine and CRF2O from https://github.com/yzhangcs/parser, StackPointer from https://github.com/XuezheMax/NeuroNLP2. The authors of TowerParse released the checkpoints trained on individual treebanks in UD2.5; as they did not provide the training script, we directly take the models trained on the largest treebank for each language, i.e., EWT for English and HDT for German.111111https://github.com/codogogo/towerparse For CoreNLP and Stanza, we use them with the out-of-the-box settings.121212We use Stanza of version 1.5.1 and CoreNLP via the client API bundled in Stanza. Note that the default models in Stanza 1.5.1 were also trained using UD2.12, however, the default German model is only trained on the GSD treebank, and the default English model is trained on the combination of 5 out of 7 train sets in the treebanks.
Except for Stackpointer, which does not have a built-in tokenizer, and CoreNLP, which has its own tokenizer, the other parsers implement the tokenizer from Stanza by default. Thus, for ease of use, we implement the Stanza tokenizer for every parser consistently, disabling sentence segmentation, as we will apply them to the curated sentences from our target corpora, for which further sentence splitting is unexpected.
@@ 4.1 Evaluation
To explore the reliability of the parsers used, we evaluate them on (1) existing UD treebanks, (2) the treebanks built upon the two target corpora of this work (denoted as “Target Treebanks”), and (3) adversarially attacked treebanks concerning two of the most severe issues identified in Section 3.
@@
We calculate the Unlabeled Attachment Score (UAS) and Labeled Attachment Score (LAS) for the parser performance. UAS and LAS are originally defined as the percentage of tokens with correctly predicted heads and head+relations respectively; in this scenario, the input of the parser is sentences pre-tokenized by humans. On the other hand, Zeman et al. (2018) test parser performance in a real-world scenario, i.e., parsing the raw text without any gold standard (e.g., sentence segmentation, tokenization, part-of-speech tags, etc.). The metrics are then re-defined as the the harmonic mean (F1) of precision P and recall R, where P is the percentage of the correct head (+relation) to the number of predicted tokens, while R is the percentage of the correct head (+relation) to the number of tokens in the gold standard.
As we will use the parsers on sentences without tokenization, we argue that the re-defined metrics on raw texts are more indicative in our case. For all evaluations, we adopt the evaluation script131313https://universaldependencies.org/conll18/evaluation.html from the CoNLL18 dependency shared task (Zeman et al., 2018).
@@ 4.1.1 Evaluation on UD Treebanks
To ensure the correctness of the training, implementation, and usage of the parsers, we first evaluate the parsers on the test sets of the existing UD treebanks, for which we choose Universal Dependencies (UD) v2.12,141414We exclude the treebank “GumReddit” for English, as there are many nonsensical annotations. and compare the results to Zeman et al. (2018). Although our evaluation setup, parsers, and the version of the UD treebanks are different from theirs, comparable (or better) results are expected since (1) the discrepancy in UD versions is small,151515Most changes are about fixing a tiny portion of incorrect annotations, e.g., see change logs on https://github.com/UniversalDependencies/UD_English-PUD/blob/master/README.md and https://github.com/UniversalDependencies/UD_German-GSD/blob/master/README.md. (2) most of the tested parsers are published later than the shared task and (3) the parsers do not need to split the text into sentences by themselves here.
@@ Results
We first compare our evaluation results on one random treebank to that reported in Zeman et al. (2018) for each language. As Figure 7 shows, the great majority of the parsers obtain higher LAS than the average LAS over parsers in Zeman et al. (2018), except for CRF2O on German, whose LAS is slightly lower than the average. Stanza, Towerparse, and Biaffine consistently outperform the best systems across different languages. Thus, we confirm the legitimacy of the used parsers.
@@
Next, we show the macro average and the standard deviation of the UAS and LAS over the UD treebanks in Table 4 (column “UD”). Biaffine consistently outperforms the others across different languages in terms of both UAS (en: 92% vs. 79%-88%; de: 87% vs. 79%-86%) and LAS (en: 89% vs. 73%-85%; de: 82% vs. 70%-80%). Furthermore, it also exhibits the strongest robustness across various treebanks, as demonstrated by the smallest average standard deviation (2.55 percentage points (pp)) over different metrics and languages. It is noteworthy that CoreNLP, the parser leveraged by all other related approaches, is the worst in our evaluation, according to all criteria.
Stanza trained on slightly mismatched data outperforms the other fine-tuned ones on average for English. Nevertheless, it is the least robust one according to its high standard deviations (6.2 pp vs. <5.5 pp in UAS; 8.0 pp vs. <6.5 pp in LAS). Interestingly, Towerparse, which was trained on mismatched data in terms of treebanks and versions, performs similarly to Biaffine on German treebanks, only 1 pp lower UAS/LAS. However, we observe that it tends to predict multiple roots for one sentence and produce cycles in the dependency tree graphs, violating the UD dependency definition. In our evaluation, it generates 440 multi-root predictions and 1,498 cycles, while the others rarely make such errors (only Stackpointer predicts multiple roots 28 times among the other parsers). Besides, it also has to skip 28 sentences due to the limitation of its implementation.161616See the authors’ explanation about skipping sentences in footnote 7 in Glavaš and Vulić (2021). CRF2O performs on par with the second-tier parsers on English data but is the worst one on German data. The reason may be that it is not capable of parsing non-projective relations, which occur more frequently in German compared to English, as shown in Gómez-Rodríguez and Ferrer-i Cancho (2017). With its built-in evaluation function, which skips the non-projective sentences in the gold standard (3k out of 23k sentences), the UAS and LAS raise from <80% to >90% on German treebanks.
@@
We illustrate the absolute difference in metrics between the sentences before and after attacking in Figure 8. Both types of attacks consistently have a greater influence on LAS compared to UAS. This can be explained by the fact that LAS is more sensitive to perturbations because it considers both attachment and labeling accuracy, whereas UAS only considers attachment accuracy. Even if the attachment remains correct, changes to the words themselves (such as spelling or character replacement) can result in incorrect labels. As depicted in Figure 7(a), historical spelling has a slight negative impact on the parsers’ performance, with a difference of less than 0.07 in LAS and that of up to around 0.05 in UAS. Among the parsers, Biaffine is the most stable one, followed by TowerParse; CoreNLP exhibits the lowest robustness.
We present the results for OCR spelling attack in Figure 7(b). OCR spelling errors have a slightly larger impact, degrading parser performance up to around 0.12 in LAS and 0.09 in UAS, when 50% characters of 2 tokens within a sentence are perturbed. Again, Biaffine is the most robust one against OCR attack, followed by CRF2O and TowerParse, which demonstrate similar sensitivity to Biaffine when the attack level is high. At the highest attack level, CoreNLP, StackPointer, and Stanza show similar sensitivity. However, at low attack levels, StackPointer and Stanza demonstrate better robustness compared to CoreNLP. In real-world cases, OCR spelling errors often involve only one character within a sentence. For example, in the sentence “… wollen Tie nun das letzte nehmen.” (from our annotation in Section 3), the only OCR error is in “Tie”, which should be corrected to “Sie”. Hence, Stanza and StackPointer are still superior to CoreNLP in real-world scenarios.
@@ 4.2 Summary
Among the explored parsers, Biaffine shows the best performance in standard evaluations as well as the highest robustness under adversarial conditions. CoreNLP and CRF2O are often the worst ones in our evaluation. Besides, they are unable to predict crossing edges, which we deem as an important indicator of language change. Non-projective sentences seem to substantially degrade CRF2O’s performance; thus, we exclude it from further experiments. CoreNLP is retained, however, for comparison purposes. Therefore, we will use five parsers to perform the large-scale analysis in Section 6: CoreNLP, StackPointer, Biaffine, TowerParse, and Stanza.
@@ 5 Metrics
is said to be the index of the main verb in a sentence. An example is shown in Figure 9, where is 6 (from root to ‘geklappt’). Lei and Jockers (2020) consider as a normalizing factor when computing the mean of dependency distances for all dependency pairs in a sentence. Liu, Zhu, and Lei (2022) showed that correlates positively with dependency distance over time in English; specifically, increases over time for long sentences while decreasing over time for short sentences.
@@
where represents the absolute difference between the position indices of a dependency pair , is a set of dependency pairs in a sentence (punctuations are excluded), length is the number of words in a sentence, and is the size . In Figure 9, mDD and nDD are 2.4 and 0.84 for example.
@@
A crossing is said to be observed when two dependency relations overlap. Ferrer-I-Cancho and Gómez-Rodríguez (2016) show that #Crossings correlates positively with dependency distance in 21 out of 30 languages, while Liu, Xu, and Liang (2017) find that machine-generated artificial texts without crossings produce shorter dependency distances compared to those with crossings. We count the number of crossings in a sentence, ignoring punctuation. An example is shown in Figure 8: the dependency relation between “Das” and “alles” crosses over the relation between “geklappt” and “hat”, i.e., .
@@
For a dependency pair, if a head follows its dependent, then the pair is termed as a head-final pair. Otherwise, if a head precedes its dependents, then the pair is considered head-initial. In linguistics, this head directionality is profiled as a crucial parameter for classifying languages (Liu, 2010). In Figure 10(a), we count the head-final cases in a sentence and also consider “head-final (%)” by computing the percentage of a head following its dependents among all pairs that contain that head. We additionally consider the average of “head-final (%)” over all the head-final pairs, termed as . We note that the sentence “Das hat alles sehr gut geklappt!” not only includes head-final dependency pairs but also head-initial pairs, e.g., the head ‘Das’ precedes its dependent ‘alles.’ Here, we consider how distant this sentence is, with a mix of head-initial and head-final pairs, compared to the same sentence but permuted to have only head-final pairs. This results in the metric that is said to be the Levenshtein edit distance (Levenshtein, 1965) between the original and permuted sentences.

### v2

@@ Results
We show the distribution of the texts identified as perfect sentences (without any issues), sentences with issues, and non-sentences in Figure 3. We observe that: (1) German data has more issues compared to English. For the older time periods, e.g., the first 3 decades in each corpus, 60-70% German sentences are found to contain issues, while there are only 20-30% English sentences found to be problematic. (2) Only a small portion of sentences were identified as non-sentences. Specifically, 6 out of 160 texts (3.75%) from DeuParl and 2 out of 220 texts (0.1%) from Hansard are found not to be sentences. (3) The proportion of problematic sentences is decreasing over time; for both languages, all sentences after 2000 are flawless.
@@
In summary, our annotations suggest that English data from Hansard is substantially less problematic than the German data from Deuparl. Notably, the most prevalent issues in the German data, namely historical spelling and OCR spelling errors, rarely appear in the English data. This leads us to question whether the Hansard Corpus has already undergone preprocessing like spelling normalization. Through contact with the UK Parliament enquiry services,101010https://www.parliament.uk/site-information/contact-us/ we received feedback indicating that the data has not been preprocessed regarding spelling normalization. As for OCR errors, a statement on their website111111https://hansard.parliament.uk/about explicitly mentions that OCR errors are curated once identified. As a consequence, we conclude that English has undergone smaller changes in terms of spelling in the past 200 years compared to German, making German the more interesting (and much less researched) language from a historical perspective.
@@ 3.2 Dataset Construction
We apply the preprocessing pipeline introduced above to get up to 200k sentences for each decade and language. To make the English and German data comparable, we limit the time period for English to the decade 1860-2020. Next, we tokenize the sentences and do part-of-speech tagging using Stanza (Qi et al., 2020). The length of sentences is then the number of tokens after tokenization. Figures 19 (in the appendix) and 5 show the distribution of sentence lengths and the average sentence length over time. We see that shorter sentences appear more often in the later periods and thus the average sentence length is overall decreasing over time for both languages. For instance, the average sentence length in DeuParl decreases from 35 to 20 in the time window 1860-1870 to 2000-2020. The only exception is the last subtrend for Hansard, where the average length increases by 2 tokens. Metrics like mDD are highly sensitive to sentence length (Lei and Jockers, 2020), thus, it is important to control for sentence length when constructing the dataset for observation.
@@ 3.2.1 Final Dataset
We select sentences with varying number of tokens for observation: 5, 10, 15, 20, 30, 40, 50, 60, and 70, covering short, middle and long sentences. We aim to observe syntactic changes in sentences of the same length (i.e., same amount of tokens). Nevertheless, to obtain sufficient sentences for each length and time period, we set an offset of 2 tokens for sentence grouping, e.g., sentences of a length from 10 to 12 are grouped together; in addition, we consider decade groups with each spanning 2 decades (except for the last one, which spans 3 decades). For both English and German, the final dataset contains 450 sentences per length and decade group. These sentences are randomly sampled from a total of 32,400 preprocessed sentences for each language.
@@ 4 Dependency Parsers
Unlike previous approaches, which solely relied on a single parser, typically the Stanford CoreNLP Parser (Manning et al., 2014), we will base our observations on various parsers with the two most popular design choices: transition-based and graph-based. Transition-based parsers sequentially predict the arc step by step based on the current state, while graph-based parsers globally optimize the dependency tree, aiming to find the highest-scored one. The parsers used in this work include: (1) transition-based: CoreNLP (Manning et al., 2014), StackPointer (Ma et al., 2018); (2) graph-based: Biaffine (Dozat and Manning, 2016), Stanza (Qi et al., 2020), TowerParse (Glavaš and Vulić, 2021) and CRF2O (Zhang, Li, and Zhang, 2020). Details of the parser configurations are provided in §9.1 in the appendix.
@@ 4.1 Evaluation
To explore the reliability of the parsers used, we evaluate them on (1) existing UD treebanks, (2) the treebanks built upon the two target corpora of this work (denoted as “Target Treebanks”), and (3) adversarially attacked treebanks concerning the most severe issues identified in §3.
@@
We calculate the Unlabeled Attachment Score (UAS) and Labeled Attachment Score (LAS) for the parser performance. UAS and LAS are originally defined as the percentage of tokens with correctly predicted heads and head+relations respectively; in this scenario, the input of the parser is sentences pre-tokenized by humans. On the other hand, Zeman et al. (2018) test parser performance in a real-world scenario, i.e., parsing the raw text without any gold standard (e.g., sentence segmentation, tokenization, part-of-speech tags, etc.). The metrics are then re-defined as the harmonic mean (F1) of precision P and recall R, where P is the percentage of the correct head (+relation) to the number of predicted tokens, while R is the percentage of the correct head (+relation) to the number of tokens in the gold standard. As we will use the parsers on sentences without tokenization, we argue that the re-defined metrics on raw texts are more indicative in our case. For all evaluations, we adopt the evaluation script from the CoNLL18 dependency shared task (Zeman et al., 2018).121212https://universaldependencies.org/conll18/evaluation.html; this script ignores the subtypes of the relations (e.g., for relation ‘acl:relcl’ (relative clause modifier), only ‘acl’ is considered.), but here, we consider the whole relations.
@@ 4.1.1 Evaluation on UD Treebanks
To ensure the correctness of the training, implementation, and usage of the parsers, we first evaluate the parsers on the test sets of the existing UD treebanks, for which we choose Universal Dependencies (UD) v2.12,131313We exclude the treebank “GUMReddit” for English, as there are many nonsensical annotations. and compare the results to Zeman et al. (2018). Although our evaluation setup, parsers, and the version of the UD treebanks are different from theirs, comparable (or better) results are expected since (1) the discrepancy in UD versions is small,141414Most changes are about fixing a tiny portion of incorrect annotations, e.g., see change logs on https://github.com/UniversalDependencies/UD_English-PUD/blob/master/README.md and https://github.com/UniversalDependencies/UD_German-GSD/blob/master/README.md. (2) most of the tested parsers are published later than the shared task and (3) the parsers do not need to split the text into sentences by themselves here.
@@ Results
In §9.2 in the appendix, we confirm the legitimacy of the used parsers by showing that the performance of our used parsers is mostly superior to those in Zeman et al. (2018). Next, we show the macro average and the standard deviation of the UAS and LAS over the UD treebanks in Table 4 (column “UD”). Biaffine consistently outperforms the others across different languages in terms of both UAS (en: 92% vs. 79%-88%; de: 87% vs. 79%-86%) and LAS (en: 89% vs. 73%-85%; de: 82% vs. 70%-80%). Furthermore, it also exhibits the strongest robustness across various treebanks, as demonstrated by the smallest average standard deviation (2.55 percentage points (pp)) over different metrics and languages. It is noteworthy that CoreNLP, the parser leveraged by all other related approaches, is the worst in our evaluation, according to all criteria.
Stanza trained on slightly mismatched data outperforms the other fine-tuned ones on average for English. Nevertheless, it is the least robust one according to its high standard deviations (6.2 pp vs. <5.5 pp in UAS; 8.0 pp vs. <6.5 pp in LAS). Interestingly, Towerparse, which was trained on mismatched data in terms of treebanks and versions, performs similarly to Biaffine on German treebanks, only 1 pp lower UAS/LAS. However, we observe that it tends to predict multiple roots for one sentence and produce cycles in the dependency tree graphs, violating the UD dependency definition. In our evaluation, it generates 440 multi-root predictions and 1,498 cycles, while the others rarely make such errors (only Stackpointer predicts multiple roots 28 times among the other parsers). Besides, it also has to skip 28 sentences due to the limitation of its implementation.151515See the authors’ explanation about skipping sentences in footnote 7 in Glavaš and Vulić (2021). CRF2O performs on par with the second-tier parsers on English data but ranks as the second-worst on German data. The reason may be that it is not capable of parsing non-projective relations, which occur more frequently in German compared to English, as shown in Gómez-Rodríguez and Ferrer-i Cancho (2017). With its built-in evaluation function, which skips the non-projective sentences in the gold standard (3k out of 23k sentences), the UAS and LAS raise from <80% to >90% on German treebanks.
@@
We illustrate the absolute difference in metrics between the sentences before and after attacking in Figure 21 in the appendix. Both types of attacks consistently have a greater influence on LAS compared to UAS. This can be explained by the fact that LAS is more sensitive to perturbations because it considers both attachment and labeling accuracy, whereas UAS only considers attachment accuracy. Even if the attachment remains correct, changes to the words themselves (such as spelling or character replacement) can result in incorrect labels. As depicted in Figure 21(a), historical spelling has a slight negative impact on the parsers’ performance, with a difference of less than 0.07 pp in LAS and that of up to around 0.05 pp in UAS. Among the parsers, Biaffine is the most stable one, followed by TowerParse; CoreNLP exhibits the lowest robustness.
We present the results for OCR spelling attack in Figure 21(b). OCR spelling errors have a slightly larger impact, degrading parser performance up to around 0.12 pp in LAS and 0.09 pp in UAS, when 50% characters of 2 tokens within a sentence are perturbed. Again, Biaffine is the most robust against OCR attack, followed by CRF2O and TowerParse, which demonstrate similar sensitivity to Biaffine when the attack level is high. At the highest attack level, CoreNLP, StackPointer, and Stanza show similar sensitivity. However, at low attack levels, StackPointer and Stanza demonstrate better robustness compared to CoreNLP. In real-world cases, OCR spelling errors often involve only one character within a sentence. For example, in the sentence “… wollen Tie nun das letzte nehmen.” (from our annotation in §3), the only OCR error is in “Tie”, which should be corrected to “Sie”. Hence, Stanza and StackPointer are still superior to CoreNLP in real-world scenarios.
@@ 4.2 Summary
All parsers exhibit robustness and show minor (or no) degradation both under adversarial attacks and the target data. Among them, Biaffine shows the best performance in standard evaluations as well as the highest robustness under adversarial conditions. CoreNLP and CRF2O are often the worst ones in our evaluation; besides, they are unable to predict crossing edges, which we deem as an important indicator of language change. Non-projective sentences seem to substantially degrade CRF2O’s performance; thus, we exclude it from further experiments. CoreNLP is retained, however, for comparison purposes. Therefore, we will use five parsers to perform the large-scale analysis in §6: CoreNLP, StackPointer, Biaffine, TowerParse, and Stanza.
@@ 5 Metrics
is the index of the word connecting to the pseudo root in a sentence. An example is shown in Figure 6, where is 6 (from root to ‘geklappt’). Lei and Jockers (2020) consider as a normalizing factor when computing the mean of dependency distances for all dependency pairs in a sentence. Liu, Zhu, and Lei (2022) showed that correlates positively with dependency distance over time in English; specifically, increases over time for long sentences while decreasing over time for short sentences.
@@
where represents the absolute difference between the position indices of a dependency pair , is a set of dependency pairs in a sentence (punctuation and root dependencies are excluded), length is the number of words in a sentence, and is the size . In Figure 6, mDD and nDD are 2.6 and 0.84 for example.
@@
A crossing is said to be observed when two dependency relations overlap. Ferrer-i Cancho and Gómez-Rodríguez (2016) show that #Crossings correlates positively with dependency distance in 21 out of 30 languages, while Liu, Xu, and Liang (2017) find that machine-generated artificial texts without crossings produce shorter dependency distances compared to those with crossings. We count the number of crossings in a sentence, ignoring punctuations. An example is shown in Figure 8: the dependency relation between “Das” and “alles” crosses over the relation between “geklappt” and “hat”, i.e., .
@@
For a dependency pair, if a head follows its dependent, then the pair is termed as a head-final pair. Otherwise, if a head precedes its dependents, then the pair is considered head-initial. In linguistics, this head directionality is profiled as a crucial parameter for classifying languages (Liu, 2010). In Figure 8(a), for each head in a sentence, we calculate the percentage of its head-final pairs to all pairs with that head (head-final(%)), and then average the head-final(%) over all heads in that sentence, termed as . We note that the sentence “Das hat alles sehr gut geklappt!” not only includes head-final dependency pairs but also head-initial pairs, e.g., the head ‘Das’ precedes its dependent ‘alles.’ Here, we consider how distant this sentence is, with a mix of head-initial and head-final pairs, compared to the same sentence but permuted to have only head-final pairs. This results in the metric that is the Levenshtein edit distance (Levenshtein, 1965) between the original and permuted sentences.